40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 3 18:48:09 2008, Alberto, Update, General, Beats of the two lasers in the absolute length measurement observed  
    Reply  Thu Jul 3 22:54:45 2008, Koji, Update, General, Beats of the two lasers in the absolute length measurement observed  
       Reply  Mon Jul 7 11:22:02 2008, Alberto, Update, General, Beats of the two lasers in the absolute length measurement observed  
          Reply  Mon Jul 7 13:06:38 2008, Koji, Update, General, Beats of the two lasers in the absolute length measurement observed  
Message ID: 634     Entry time: Thu Jul 3 18:48:09 2008     Reply to this: 635
Author: Alberto 
Type: Update 
Category: General 
Subject: Beats of the two lasers in the absolute length measurement observed  
I adjusted the alignment of the flipper mirror as suggested by Koji making the two beam spots match. I also aligned all the IFO mirrors (ITMs, PRM, SRM, ETMs) to have more power for the IFO signal at the AS port. When I did that I could see the beats at the AS OSA. Then I explored the range of temperature of the NPRO from 35deg (C) to 51.2807deg and at that point I could observe a peak corresponding to the beat at about 10MHz on the network analyzer. The peak tends to drift because the laser takes probably a longer time to actually thermalize and it moves very rapidly changing the temperature of the laser.
ELOG V3.1.3-